OpenStudy (anonymous):

∬_D y d A D is between x=y^2 -1 & x=-y^2+2y+3

OpenStudy (zarkon):

\[\int\limits_{-1}^{2}\int\limits_{y^2-1}^{-y^2+2y+3}y\,dx\,dy=\frac{9}{2}\]

OpenStudy (anonymous):

Zarkon, shouldn't the outter integral be: \[\huge\int\limits_0^3\] Instead of: \[\huge\int\limits_0^2\]

OpenStudy (zarkon):

solve y^2 -1=-y^2+2y+3 you get y=-1 and 2

OpenStudy (anonymous):

Ahhhh gotcha lol